The atomic mass of 5626Fe is 55.934939 u, and the atomic mass of 5627Co is 55.939847 u.

How much kinetic energy will the products of the decay have?

Express your answer with the appropriate units.
Answer #1

Write the expression to calculate the mass defect. Am- m(Co-m(Fe) (55.939847 u)-(55.934939 u) 4908x10% u Calculate the bindinCalculate the energy carried by the positron and the electron = 2(9.11×10-31 kg)(3x108 m/s*) = 1.63×10-15 J Convert the above